OpenStudy (anonymous):

I have one more chance to get this right... 60° 50' 40" - 30° 40' 50" _______________ __° __' __" ???

OpenStudy (anonymous):

\[30^{0}9'50"\]

OpenStudy (anonymous):

(60-50)+40 =50 49-40=9 60-30=30